

********************************************************************
UARS WEEKLY STATUS REPORT - 11 MAR 2005
********************************************************************
***********
SOLSTICE
***********
This week, SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
46% of each day's data amounting to about 524 minutes of solar
observing time per day.
***********
SUSIM
***********
During this past week, SUSIM successfully gathered daily mid (1.1 nm)
resolution, MgII, occultation, solar pointing offset, and continuous
monitoring scans for time periods with UARS telemetry or tape
recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004299 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004300 and 2004359 are currently waiting archival at GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004309. MEPS data between 2004310 and 2004329
are currently waiting archival at GSFC DAAC.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE returns to successful operations after the graze period. All
engineering parameters are within nominal operating values and the
exoatmospheric HF science channel voltage is nominal. All channels are
still being evaluated. The UARS FOT continues to provide exceptional
help and attention during the HALOE power on. The power on process was
executed in less than twelve minutes, allowing HALOE to capture an
immediate Sunset.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 26 February 2005 (Orbit 73677, GMT Day 057) through
04 March 2005 (Orbit 73781, GMT Day 063).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 058/1353 73701 DAY ONLY TO FULLTIME
PEM 058/1432 73701 AXIS 1 & 2 ON
HRDI 062/1935 73764 RESTART
The beta angle ranged between the angles of 52.2 degrees to 63.2
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 3.0 degrees on 21 March 2005 (DOY 080).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 3.91 to 3.23 degrees C. Additional
battery performance data for each battery on 25 February 2005 and 04
March 2005 are provided in Appendix D.
The clock error ranged from 10.1 to -09.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 02
HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to monitor the batteries closely. The MPS
continues to utilize the VT-7 charge rate.
The FOT has completed their CMS software acceptance testing. All
tests results were positive and the official release of the new
software was installed on the backup CMS. With the install complete,
the FOT will run a few simulated days of operations to confirm the
software. Phase 1 of the simulations has been completed. Some
additional testing/verification will be required next week, as well as
uplink tests to the spacecraft. If all goes well, then the new
release will be put onto the operational CMS.
MEETINGS
The FOT attended the UARS monthly status review on 01 March 2005 to
discuss the significant events for the month of February.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 21 March 2005 (DOY 080). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 04 March 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
73701 058/1353 05-054 PEM PEM FROM DAY ONLY TO FULLTIME
73701 058/1432 05-055 PEM AXIS 1 & 2 ON
73707 059/0008 05-056 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73748 061/1723 05-057 OBC DUMP OBC BANKS
73764 062/1935 05-059 HDRI RESTART (BEGIN)
73765 062/2118 05-059 HRDI RESTART (COMPLETE)
73766 062/2257 05-058 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 04 March 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 04 March 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
73707 059/0033 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
73766 062/2308 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 04 March 2005
25 February 2005 - GMT Day 056 Beta = 49.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -1.79/-3.05 +00.0/+00.0 07.0 0.00 0.0
2 168.0/ 56.0 +10.40/+9.06 +26.4/-15.2 26.7 1.11 17.0
3 11.2/- 89.6 +7.09/+5.15 +22.4/-14.4 26.7 1.07 16.0
04 March 2005 - GMT Day 063 Beta = 63.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 201.6/ 89.6 +9.06/+7.74 +20.8/-14.4 27.2 1.10 16.8
3 -22.4/-100.8 +5.15/+4.51 +18.8/-13.2 27.0 1.06 15.9
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 04 March 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-011 FOT 73699 058/1213 38687 Negative Acquisition
05-012 FOT 73769 063/0306 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-010 FOT 73616 052/2218 38591 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-005 FOT See List Feb 05 30832 Generic Late Acquisition
05-002 MPS 73008 012/1251 N/A BATT 3 Diff Voltage High
05-011 FOT 73699 058/1213 38687 Negative Acquisition
05-012 FOT 73769 063/0306 30832 Generic Late Acquisition